# Universality of Efimov states in highly mass-imbalanced cold-atom mixtures with van der Waals and dipole interactions

### v2

@@ Abstract
We study three-body systems in a mass-imbalanced two-component cold-atom mixture, and investigate the three-body parameter of their Efimov states for both bosonic and fermionic systems, with a major focus on the Er-Er-Li Efimov states. For a system interacting solely via van der Waals interactions, the van der Waals universality of the three-body parameter is analytically derived using the quantum defect theory. With the addition of a perturbative dipole interaction between the heavy atoms, the three-body parameters of the bosonic and fermionic Efimov states are found to behave differently. When the dipole interaction is as strong as the van der Waals interaction, corresponding to realistic Er-Er-Li Efimov states, we show that the van der Waals universality persists once the effects of the non-perturbative dipole interaction are renormalized into the $s$-wave and $p$-wave scattering parameters between the heavy atoms. For a dipole interaction much stronger than the van der Waals interaction, we find that the universality of the Efimov states can be alternatively characterized by a quasi-one-dimensional scattering parameter due to a strong anisotropic deformation of the Efimov wavefunctions. Our work thus clarifies the interplay of isotropic and anisotropic forces in the universality of the Efimov states. Based on the renormalized van der Waals universality, the three-body parameter is estimated for specific isotopes of Er-Li cold-atom mixtures.
@@ I Introduction
Recent progress in the field of cold atoms has significantly advanced our understanding of strongly correlated quantum systems and their universal behaviors. A prime example of quantum phenomenon realized in cold atoms is the Efimov effect [1, 2, 3, 4, 5, 6], which allows the existence of weakly bound three-body states featuring a discrete scale invariance. While these Efimov states have been theoretically suggested to appear universally for various physical systems [3, 7, 8], they require the particles to interact with large $s$-wave scattering lengths. This condition can be achieved in cold atoms by using a Feshbach resonance [9, 10], which enables the precise control of the $s$-wave scattering length between the atoms. Efimov states have been observed in systems of identical bosons [11], mass-imbalanced bosonic mixtures [12, 13, 14], and three-component Fermi systems [15, 16, 17]. An Efimov state has also been observed as the excited state of the 4He trimer [18], thanks to a naturally large $s$-wave scattering length between 4He atoms. Importantly, the Efimov states observed for various different atomic species [19] and internal spin states [20] have been found to have almost the same physical property: the three-body parameter, characterizing the binding energy and hence the size of the Efimov states, has been found to scale universally with the van der Waals length between the atoms. This universality of the three-body parameter has been termed “van der Waals universality” and actively investigated both theoretically and experimentally for identical bosons [19, 20, 21, 22, 23] as well as mass-imbalanced bosonic mixtures [24, 25]; It has been demonstrated to hold true except for systems close to a narrow Feshbach resonance [26, 27, 28, 29, 30, 31, 32, 33, 34].
@@
Efimov physics has also been explored in fermionic systems. The repulsion between identical fermions due to the Pauli exclusion competes with the Efimov attraction, and leads to a richer few-body physics. Theoretically, Efimov trimers have been predicted to appear in a mass-imbalanced two-component fermionic mixture when the mass ratio is large: $M/m>13.6\ldots$ [2, 35]. With such a large mass imbalance, the Efimov attraction surpasses the repulsion and forms the Efimov trimers, in stark contrast to a moderate mass ratio $8.1\ldots<M/m<13.6\ldots$ for which there exist “Kartavtsev-Malykh trimers” [36, 37] and their crossover into the Efimov trimers [38], and a small mass ratio $M/m<8.1\ldots$ for which no trimer appears. While these fermionic trimers are formed in the $L^{\Pi}=1^{-}$ state, distinct from the $L=0^{+}$ state of identical bosons, they require a fermionic cold atom mixture with a well-controlled Feshbach resonance. Until recently, the 40K-6Li mixture has been the major mass-imbalanced Fermi system. The mass ratio, however, is smaller than that required for the appearance of the trimer, and only a precursor of the trimer formation has been observed [39]. Furthermore, the Feshbach resonance of 40K-6Li is of a relatively narrow character [40, 10]. Recently, Feshbach resonances have been realized in highly mass-imbalanced Fermi mixtures of Er-Li [41, 42] and Cr-Li [43], where multiple broad resonances have been found. This opens the possibility of observing the trimer physics of fermions in cold atoms. In particular, Er-Li is well inside the Efimov regime, and it is expected to be the first system to exhibit a “rotating” Efimov state, i.e. Efimov state with non-zero orbital angular momentum $L=1$. One major challenge of the Er-Li mixture is that there is a strong dipole-dipole interaction between the Er atoms [44, 45]. The dipole interaction couples states with different angular momentum, so that the orbital angular momentum $L$ is no longer a good quantum number [46, 47]. Since the dipole interaction is as strong as the van der Waals interaction, we are faced with the question of how the interplay of dipole and van der Waals interaction can modify the universal properties of the Efimov states: “Is the three-body parameter in the Er-Li system universal?”
@@
We address this question by investigating the Efimov states in a highly mass-imbalanced three-body system. Using the Born-Oppenheimer approximation valid for a highly mass-imbalanced system, we study the universality of the three-body parameter, i.e. the binding energy of the Efimov states in the unitary limit. To elucidate the role of the statistics of the heavy atoms, we study both bosonic and fermionic systems with a variable strength of dipole interaction. First, we study the three-body system interacting solely with a van der Waals interaction. Using the quantum defect theory (QDT), the three-body parameter is explicitly expressed as a universal function of the van der Waals length and $s$-wave scattering length ($p$-wave scattering volume) between the heavy atoms for bosons (fermions), respectively. Second, for a system interacting with the van der Waals interaction and a weak dipole interaction, our perturbative analysis, together with numerical calculations, show that the bosonic and fermonic Efimov states are affected in a distinct manner by the dipole interaction. Third, for moderate dipole interaction strengths, including the realistic dipole strengths between Er atoms, our coupled-channel numerical calculations vindicate the universality of the three-body parameter. For the bosonic system, the van der Waals universality formula is found to persist even though the dipole interaction is so large that the $s$-wave scattering length and the binding energy are significantly modified; once the effects of the dipole interaction are renormalized into the $s$-wave scattering length, the QDT formula derived for a purely van der Waals system holds true, so that the $s$-wave scattering length between the heavy atoms universally characterize the three-body parameter. Based on this renormalized van der Waals universality, we estimate the three-body parameters of the Efimov states for some specific isotopes of Er-Li systems. For the fermionic systems, the $p$-wave scattering volume is no longer a viable physical parameter, and we find an alternative low-energy scattering parameter which characterizes the universality of the three-body parameter. Finally, for a dipole interaction much stronger than the van der Waals interaction, the above universal description with the three-dimensional scattering parameters should break down, but we find that quasi-one-dimensional scattering parameters between the heavy atoms can be used to universally describe the Efimov states in both bosonic and fermionic systems.
@@
This paper is organized as follows: in Sec. II, after introducing the Born-Oppenheimer description of the mass-imbalanced three-body system, our QDT analysis of non-dipolar system is presented. In Sec. III, the results of our numerical calculations are presented and compared with our analytical results for non-dipolar, weak dipole, and moderate dipole interactions, followed by the analysis in the limit of strong dipole interactions. In Sec. IV, we present our estimates of the three-body parameters of some Er-Li isotopes. We conclude and present future perspectives in Sec. V. Throughout the paper, the natural unit $\hbar=1$ is used.
@@ II.1 General Formulation
In a highly mass-imbalanced system, the Born-Oppenheimer approximation provides a good description of the Efimov physics. Indeed, it successfully reproduces the Efimov scale factor, and the critical mass ratio for the Efimov effect $M/m=13.99\ldots$ [2] in good agreement with the exact value $M/m=13.60\ldots$ [35]. The Born-Oppenheimer approximation has also been confirmed to reproduce well the van der Waals universality of the Efimov states without the dipole interaction for mass ratios $14\lesssim M/m\lesssim 29$ [24]. It is therefore a good approximation for an Er-Li mixture whose mass ratio is $M/m\approx 28$. The light particle’s Schrödinger equation is solved analytically and leads to the induced attractive interaction between the heavy particles $-\Omega^{2}/2mr^{2}$, where $\Omega=0.5671\ldots$ [3]. The Schrödinger equation between the heavy particles then reads
@@
where $\theta$ is an angle measured from the $z$-axis taken to be parallel to the orientation of the dipoles (parallel to the external magnetic field). Here, $C_{6}$ and $C_{dd}$ are the van der Waals and dipole coefficients, from which we can define the van der Waals length $r_{\mathrm{vdw}}=\dfrac{1}{2}(MC_{6})^{1/4}$ and the dipole length $a_{dd}=MC_{dd}/3$ between the heavy atoms.
The binding energy of the Efimov states (i.e. three-body parameter) can be obtained by solving Eq. (1). It needs to be supplemented with a short-range boundary condition. For this purpose, we use a hard wall condition at $R_{\mathrm{min}}$: $\psi(r=R_{\mathrm{min}})=0$. We will see later that the results are insensitive to the way the boundary condition is introduced as long as $R_{\mathrm{min}}$ is small enough. As the dipole interaction couples different partial waves, the angular momentum is no longer a good quantum number, but the parity $\Pi$ and the $z$-component of the angular momentum $M_{z}$ are good quantum numbers. While various values of $M_{z}$ are allowed in general, we restrict ourselves to the $M_{z}=0$ channel in this paper. For the bosonic heavy atoms, the $M_{z}^{\Pi}=0^{+}$ state is allowed by the symmetry. This state involves the $L=0,2,4,...$ angular momentum channels, among which $L^{\Pi}=0^{+}$ channel shows the Efimov effect and the others do not for the Er-Li mass ratio [2, 36, 49, 50, 3]. For the fermionic heavy atoms, $M_{z}^{\Pi}=0^{-}$ channel is allowed by the symmetry. This state involves $L=1,3,5,...$ angular momentum channels, among which the $L^{\Pi}=1^{-}$ channel shows the Efimov effect and the others do not for the mass ratio in question [2, 36, 49, 50, 3]. Therefore, in the following, we study the $M_{z}^{\Pi}=0^{+}$ and $M_{z}^{\Pi}=0^{-}$ states for the bosonic and fermionic Er atoms, respectively.
@@ II.2 Quantum Defect Theory for non-dipolar system
with a short-range boundary condition $\psi(|\bm{r}|=R_{\mathrm{min}})=0$. While this equation was studied numerically in Ref. [25] and the van der Waals universality of the Efimov states was clarified for the mass-imbalanced system, we show below that Eq. (2) can be solved analytically and provide explicit formulae for the binding energies and wavefunctions of the Efimov states.
@@
Since we are interested in the Efimov states, we focus on the $\ell=0$ ($\ell=1$) channel where $s_{\ell}^{2}<0$ for the mass ratio of interests $M/m\approx 28$ for bosons (fermions), respectively [2, 36, 3, 49, 50]. Equation (3) can be solved analytically using Ref. [51], with a remark that the centrifugal term is attractive due to the Efimov attraction. In other words, the angular momentum in Ref. [51] should be generalized to an imaginary number. We can then apply the quantum defect theory (QDT) [52, 53] to obtain the analytical behavior of the Efimov states; using the QDT parameter $K^{c}$ introduced as
@@
$K^{c}=\frac{W^{c}_{f-}}{W^{c}_{g-}}=\frac{(1+M_{\ell})\sin\dfrac{\pi\nu}{2}\cos\theta_{\ell}+(1-M_{\ell})\cos\dfrac{\pi\nu}{2}\sin\theta_{\ell}}{(1-M_{\ell})\cos\dfrac{\pi\nu}{2}\cos\theta_{\ell}+(1+M_{\ell})\sin\dfrac{\pi\nu}{2}\sin\theta_{\ell}}.$ (6)
@@
the only difference with Eq. (2) is the presence/absence of the $1/r^{2}$ attraction. At short distance $r\sim R_{\mathrm{min}}\ll r_{\mathrm{vdw}}$, this attraction is much smaller than the van der Waals force so that the two equations are essentially the same. More physically, the short-range phase of the three-body system is dominated by the potential between the heavy atoms. This assumption should be particularly true for a highly mass-imbalanced system where the atomic polarizability of the heavy atoms, hence the van der Waals force and higher-order forces are much stronger than those for the light atom. Therefore, the QDT parameter of the three-body system can be approximated by that of two heavy atoms $K^{c}_{\mathrm{3body}}\simeq K^{c}_{\mathrm{2body}}$, which amounts to adopting the same short-range cutoff $R_{\mathrm{min}}$ value for the two- and three-body systems. While this assumption is plausible for our model, it can be compromised by finite-range effects of the heavy-light atoms’ interaction. It may also be invalidated by the non-adiabatic effects beyond the Born-Oppenheimer approximation, which break down the single-channel description.
@@
where $a^{(\mathrm{HH})}$ and $v_{p}^{(\mathrm{HH})}$ are $s$-wave scattering length and $p$-wave scattering volume between the heavy atoms, respectively. Noting that the bosonic (fermionic) heavy atoms interact dominantly with the $s$-wave scattering length ($p$-wave scattering volume) at low energy, and substituting Eqs. (13) and (14) into Eq. (7), the binding energies of the Efimov states are represented as
@@
As we will see in Sec. III.1, these analytical formulae excellently reproduce the numerical results for non-dipolar systems. These analytical formulae suggest that the binding energy of the Efimov states in the unitary limit, i.e. the three-body parameter, is insensitive to the short-range details of the system but universally determined by the van der Waals length and the $s$-wave scattering length ($p$-wave scattering volume) for the bosonic (fermionic) Efimov states. This van der Waals universality has been experimentally demonstrated for equal-mass [19, 20] and mass-imbalanced bosonic systems [29] close to a broad Feshbach resonance. The van der Waals universality has also been demonstrated theoretically with numerical calculations for equal-mass [23, 21, 22] and mass-imbalanced bosonic systems [24, 25]. In particular, the same Born-Oppenheimer equation (2) with the same short-range boundary condition has been employed in Ref. [25] and solved numerically to study the universality of the Efimov states. Our QDT analysis provides an analytical demonstration of the van der Waals universality of hetero-nuclear Efimov states; namely, the assumption of $K^{c}_{\mathrm{3body}}\simeq K^{c}_{\mathrm{2body}}$ is pivotal in universally characterizing the three-body parameter with the two-body scattering parameters between the heavy atoms. As we use the single-channel QDT, Eqs. (15) and (16) are invalid for describing the Efimov states near a narrow Feshbach resonance, which have been demonstrated both experimentally [26, 27, 28, 29] and theoretically [30, 31, 32, 33, 34] to deviate from the van der Waals universality; the multi-channel QDT [55, 56] is necessary for capturing their sensitivity to the channel coupling and nature of the internal spin states.

### v3

@@ Abstract
We study three-body systems in a mass-imbalanced two-component cold-atom mixture, and we investigate the three-body parameter of their Efimov states for both bosonic and fermionic systems, with a major focus on the Er-Er-Li Efimov states. For a system interacting solely via van der Waals interactions, the van der Waals universality of the three-body parameter is analytically derived using the quantum defect theory. With the addition of a perturbative dipole interaction between the heavy atoms, the three-body parameters of the bosonic and fermionic Efimov states are found to behave differently. When the dipole interaction is as strong as the van der Waals interaction, corresponding to realistic Er-Er-Li Efimov states, we show that the van der Waals universality persists once the effects of the non-perturbative dipole interaction are renormalized into the $s$-wave and $p$-wave scattering parameters between the heavy atoms. For a dipole interaction much stronger than the van der Waals interaction, we find that the universality of the Efimov states can be alternatively characterized by a quasi-one-dimensional scattering parameter due to a strong anisotropic deformation of the Efimov wave functions. Our work thus clarifies the interplay of isotropic and anisotropic forces in the universality of the Efimov states. Based on the renormalized van der Waals universality, the three-body parameter is estimated for specific isotopes of Er-Li cold-atom mixtures.
@@ I Introduction
Recent progress in the field of cold atoms has significantly advanced our understanding of strongly correlated quantum systems and their universal behaviors. A prime example of quantum phenomenon realized in cold atoms is the Efimov effect [1, 2, 3, 4, 5, 6], which allows the existence of weakly bound three-body states featuring a discrete scale invariance. While these Efimov states have been theoretically suggested to appear universally for various physical systems [3, 7, 8], they require the particles to interact with large $s$-wave scattering lengths. This condition can be achieved in cold atoms by using a Feshbach resonance [9, 10], which enables the precise control of the $s$-wave scattering length between the atoms. Efimov states have been observed in systems of identical bosons [11], mass-imbalanced bosonic mixtures [12, 13, 14], and three-component Fermi systems [15, 16, 17]. An Efimov state has also been observed as the excited state of the 4He trimer [18], due to a naturally large $s$-wave scattering length between 4He atoms. Importantly, the Efimov states observed for various different atomic species [19] and internal spin states [20] have been found to have almost the same physical property: the three-body parameter, characterizing the binding energy and hence the size of the Efimov states, has been found to scale universally with the van der Waals length between the atoms. This universality of the three-body parameter has been termed “van der Waals universality” and actively investigated both theoretically and experimentally for identical bosons [19, 20, 21, 22, 23] as well as mass-imbalanced bosonic mixtures [24, 25]; It has been demonstrated to hold true except for systems close to a narrow Feshbach resonance [26, 27, 28, 29, 30, 31, 32, 33, 34].
@@
Efimov physics has also been explored in fermionic systems. The repulsion between identical fermions due to the Pauli exclusion competes with the Efimov attraction, and leads to a richer few-body physics. Theoretically, Efimov trimers have been predicted to appear in a mass-imbalanced two-component fermionic mixture when the mass ratio is large: $M/m>13.6\ldots$ [2, 35]. With such a large mass imbalance, the Efimov attraction surpasses the repulsion and forms the Efimov trimers, in stark contrast to a moderate mass ratio $8.1\ldots<M/m<13.6\ldots$ for which there exist “Kartavtsev-Malykh trimers” [36, 37] and their crossover into the Efimov trimers [38], and a small mass ratio $M/m<8.1\ldots$ for which no trimer appears. While these fermionic trimers are formed in the $L^{\Pi}=1^{-}$ state, distinct from the $L=0^{+}$ state of identical bosons, they require a fermionic cold atom mixture with a well-controlled Feshbach resonance. Until recently, the 40K-6Li mixture has been the major mass-imbalanced Fermi system. The mass ratio, however, is smaller than that required for the appearance of the trimer, and only a precursor of the trimer formation has been observed [39]. Furthermore, the Feshbach resonance of 40K-6Li is of a relatively narrow character [40, 10]. Recently, Feshbach resonances have been realized in highly mass-imbalanced Fermi mixtures of Er-Li [41, 42] and Cr-Li [43], where multiple broad resonances have been found. This opens the possibility of observing the trimer physics of fermions in cold atoms. In particular, Er-Li is well inside the Efimov regime, and it is expected to be the first system to exhibit a “rotating” Efimov state, i.e., Efimov state with non-zero orbital angular momentum $L=1$. One major challenge of the Er-Li mixture is that there is a strong dipole-dipole interaction between the Er atoms [44, 45]. The dipole interaction couples states with different angular momentum, so that the orbital angular momentum $L$ is no longer a good quantum number [46, 47]. Since the dipole interaction is as strong as the van der Waals interaction, we are faced with the question of how the interplay of dipole and van der Waals interaction can modify the universal properties of the Efimov states: “Is the three-body parameter in the Er-Li system universal?”
@@
We address this question by investigating the Efimov states in a highly mass-imbalanced three-body system. Using the Born-Oppenheimer approximation valid for a highly mass-imbalanced system, we study the universality of the three-body parameter, i.e., the binding energy of the Efimov states in the unitary limit. To elucidate the role of the statistics of the heavy atoms, we study both bosonic and fermionic systems with a variable strength of dipole interaction. First, we study the three-body system interacting solely with a van der Waals interaction. Using the quantum defect theory (QDT), the three-body parameter is explicitly expressed as a universal function of the van der Waals length and $s$-wave scattering length ($p$-wave scattering volume) between the heavy atoms for bosons (fermions), respectively. Second, for a system interacting with the van der Waals interaction and a weak dipole interaction, our perturbative analysis, together with numerical calculations, show that the bosonic and fermonic Efimov states are affected in a distinct manner by the dipole interaction. Third, for moderate dipole interaction strengths, including the realistic dipole strengths between Er atoms, our coupled-channel numerical calculations vindicate the universality of the three-body parameter. For the bosonic system, the van der Waals universality formula is found to persist even though the dipole interaction is so large that the $s$-wave scattering length and the binding energy are significantly modified; once the effects of the dipole interaction are renormalized into the $s$-wave scattering length, the QDT formula derived for a purely van der Waals system holds true, so that the $s$-wave scattering length between the heavy atoms universally characterizes the three-body parameter. Based on this renormalized van der Waals universality, we estimate the three-body parameters of the Efimov states for some specific isotopes of Er-Li systems. For the fermionic systems, the $p$-wave scattering volume is no longer a viable physical parameter, and we find an alternative low-energy scattering parameter which characterizes the universality of the three-body parameter. Finally, for a dipole interaction much stronger than the van der Waals interaction, the above universal description with the three-dimensional scattering parameters should break down, but we find that quasi-one-dimensional scattering parameters between the heavy atoms can be used to universally describe the Efimov states in both bosonic and fermionic systems.
@@
This paper is organized as follows: in Sec. II, after introducing the Born-Oppenheimer description of the mass-imbalanced three-body system, our QDT analysis of non-dipolar system is presented. In Sec. III, we present the results of our numerical calculations and compare them with our analytical results for non-dipolar, weak dipole, and moderate dipole interactions, followed by the analysis in the limit of strong dipole interactions. In Sec. IV, we present our estimates of the three-body parameters of some Er-Li isotopes. We conclude and present future perspectives in Sec. V. Throughout the paper, the natural unit $\hbar=1$ is used.
@@ II.1 General Formulation
In a highly mass-imbalanced system, the Born-Oppenheimer approximation provides a good description of the Efimov physics. Indeed, it successfully reproduces the Efimov scale factor, and the critical mass ratio for the Efimov effect $M/m=13.99\ldots$ [2] in good agreement with the exact value $M/m=13.60\ldots$ [35]. The Born-Oppenheimer approximation has also been confirmed to reproduce well the van der Waals universality of the Efimov states without the dipole interaction for mass ratios $14\lesssim M/m\lesssim 29$ [24]. It is therefore a good approximation for an Er-Li mixture, whose mass ratio is $M/m\approx 28$. The light particle’s Schrödinger equation is solved analytically and leads to the induced attractive interaction between the heavy particles $-\Omega^{2}/2mr^{2}$, where $\Omega=0.5671\ldots$ [3]. The Schrödinger equation between the heavy particles then reads
@@
where $\theta$ is an angle measured from the $z$ axis taken to be parallel to the orientation of the dipoles (parallel to the external magnetic field). Here, $C_{6}$ and $C_{dd}$ are the van der Waals and dipole coefficients, from which we can define the van der Waals length $r_{\mathrm{vdw}}=\dfrac{1}{2}(MC_{6})^{1/4}$ and the dipole length $a_{dd}=MC_{dd}/3$ between the heavy atoms.
The binding energy of the Efimov states (i.e., three-body parameter) can be obtained by solving Eq. (1). It needs to be supplemented with a short-range boundary condition. For this purpose, we use a hard wall condition at $R_{\mathrm{min}}$: $\psi(r=R_{\mathrm{min}})=0$. We will see later that the results are insensitive to the way the boundary condition is introduced as long as $R_{\mathrm{min}}$ is small enough. As the dipole interaction couples different partial waves, the angular momentum is no longer a good quantum number, but the parity $\Pi$ and the $z$ component of the angular momentum $M_{z}$ are good quantum numbers. While various values of $M_{z}$ are allowed in general, we restrict ourselves to the $M_{z}=0$ channel in this paper. For the bosonic heavy atoms, the $M_{z}^{\Pi}=0^{+}$ state is allowed by the symmetry. This state involves the $L=0,2,4,...$ angular momentum channels, among which the $L^{\Pi}=0^{+}$ channel shows the Efimov effect and the others do not for the Er-Li mass ratio [2, 36, 49, 50, 3]. For the fermionic heavy atoms, the $M_{z}^{\Pi}=0^{-}$ channel is allowed by the symmetry. This state involves $L=1,3,5,...$ angular momentum channels, among which the $L^{\Pi}=1^{-}$ channel shows the Efimov effect and the others do not for the mass ratio in question [2, 36, 49, 50, 3]. Therefore, in the following, we study the $M_{z}^{\Pi}=0^{+}$ and $M_{z}^{\Pi}=0^{-}$ states for the bosonic and fermionic Er atoms, respectively.
@@ II.2 Quantum Defect Theory for non-dipolar system
with a short-range boundary condition $\psi(|\bm{r}|=R_{\mathrm{min}})=0$. While this equation was studied numerically in Ref. [25] and the van der Waals universality of the Efimov states was clarified for the mass-imbalanced system, we show below that Eq. (2) can be solved analytically and provide explicit formulas for the binding energies and wave functions of the Efimov states.
@@
Since we are interested in the Efimov states, we focus on the $\ell=0$ ($\ell=1$) channel where $s_{\ell}^{2}<0$ for the mass ratio of interest $M/m\approx 28$ for bosons (fermions), respectively [2, 36, 3, 49, 50]. Equation (3) can be solved analytically using Ref. [51], with a remark that the centrifugal term is attractive due to the Efimov attraction. In other words, the angular momentum in Ref. [51] should be generalized to an imaginary number. We can then apply the quantum defect theory (QDT) [52, 53] to obtain the analytical behavior of the Efimov states; using the QDT parameter $K^{c}$ introduced as
@@
$K^{c}=\frac{W^{c}_{f-}}{W^{c}_{g-}}=\frac{(1+M_{\ell})\sin\dfrac{\pi\nu}{2}\cos\theta_{\ell}+(1-M_{\ell})\cos\dfrac{\pi\nu}{2}\sin\theta_{\ell}}{(1-M_{\ell})\cos\dfrac{\pi\nu}{2}\cos\theta_{\ell}-(1+M_{\ell})\sin\dfrac{\pi\nu}{2}\sin\theta_{\ell}}.$ (6)
@@
the only difference with Eq. (2) is the presence/absence of the $1/r^{2}$ attraction. At short distance $r\sim R_{\mathrm{min}}\ll r_{\mathrm{vdw}}$, this attraction is much smaller than the van der Waals force so that the two equations are essentially the same. More physically, the short-range phase of the three-body system is dominated by the potential between the heavy atoms. This assumption should be particularly true for a highly mass-imbalanced system, because due to larger atomic polarizability of the heavy atoms, their van der Waals force and higher-order forces are much stronger than those for the light atom. Therefore, the QDT parameter of the three-body system can be approximated by that of two heavy atoms $K^{c}_{\mathrm{3body}}\simeq K^{c}_{\mathrm{2body}}$, which amounts to adopting the same short-range cutoff $R_{\mathrm{min}}$ value for the two- and three-body systems. While this assumption is plausible for our model, it can be compromised by finite-range effects of the heavy-light atoms’ interaction. It may also be invalidated by the non-adiabatic effects beyond the Born-Oppenheimer approximation, which break down the single-channel description.
@@
where $a^{(\mathrm{HH})}$ and $v_{p}^{(\mathrm{HH})}$ are the $s$-wave scattering length and $p$-wave scattering volume between the heavy atoms, respectively. Noting that the bosonic (fermionic) heavy atoms interact dominantly with the $s$-wave scattering length ($p$-wave scattering volume) at low energy, and substituting Eqs. (13) and (14) into Eq. (7), the binding energies of the Efimov states are represented as
@@
As we will see in Sec. III.1, these analytical formulas excellently reproduce the numerical results for non-dipolar systems. These analytical formulas suggest that the binding energy of the Efimov states in the unitary limit, i.e. the three-body parameter, is insensitive to the short-range details of the system but universally determined by the van der Waals length and the $s$-wave scattering length ($p$-wave scattering volume) for the bosonic (fermionic) Efimov states. This van der Waals universality has been experimentally demonstrated for equal-mass [19, 20] and mass-imbalanced bosonic systems [29] close to a broad Feshbach resonance. The van der Waals universality has also been demonstrated theoretically with numerical calculations for equal-mass [23, 21, 22] and mass-imbalanced bosonic systems [24, 25]. In particular, the same Born-Oppenheimer equation (2) with the same short-range boundary condition has been employed in Ref. [25] and solved numerically to study the universality of the Efimov states. Our QDT analysis provides an analytical demonstration of the van der Waals universality of hetero-nuclear Efimov states; namely, the assumption of $K^{c}_{\mathrm{3body}}\simeq K^{c}_{\mathrm{2body}}$ is pivotal in universally characterizing the three-body parameter with the two-body scattering parameters between the heavy atoms. As we use the single-channel QDT, Eqs. (15) and (16) are invalid for describing the Efimov states near a narrow Feshbach resonance, which have been demonstrated both experimentally [26, 27, 28, 29] and theoretically [30, 31, 32, 33, 34] to deviate from the van der Waals universality; the multi-channel QDT [55, 56] is necessary for capturing their sensitivity to the channel coupling and nature of the internal spin states.